# Revisiting the Equivalence of Bayesian Neural Networks and Gaussian Processes: On the Importance of Learning Activations

### v2

@@ 1 Introduction
Function-space priors for BNNs offer a better way of specifying beliefs on data modeled by the model. Unlike priors on model parameters (i.e. weights), which lack interpretability and are hard to specify, they ultimately lead to more intuitive and meaningful representation of prior knowledge (Sun et al., , 2019). Function-space priors can be conveniently specified in terms of GPs. Instead of specifying complex distributions over model parameters, GPs allow for defining priors over entire functions, through the choice of kernel capturing requirements such as smoothness or periodicity.
@@
Finding a GP that matches a wide BNN is straightforward. A classic result by Neal, (1996); Williams, (1996), later extended to deep neural networks by Lee et al., (2017); Matthews et al., (2018), shows that wide layers in neural networks behave a priori like GPs. This is achieved by identifying GP kernels matching the covariances of (B)NNs.
@@
In contrast, finding BNNs that exhibit the desired behavior of GPs is a notoriously difficult problem, with solutions or approximations existing for only a few GP kernels. For example, Meronen et al., (2020) recently found a solution for the popular Matérn kernel using the Wiener-Khinchin theorem. However, the proposed BNN activation assumes binary white noise priors on the model weights, which severely hinders posterior learning. Consequently, an approximate solution with Gaussian priors must be used, ultimately leading to suboptimal performance, as we demonstrate in Sec. 4.3.
@@
The challenge of imposing function-space a priori behavior on BNNs previously has been addressed using gradient-based optimization. In particular, Flam-Shepherd et al., (2017), Flam-Shepherd et al., (2018), and especially Tran et al., (2022) attempted to solve this by learning priors on parameters. However, to achieve sufficient fidelity with simple priors (Gaussian or hierarchical), their approach requires deep networks, which presents a challenge for posterior learning due to the lack of effective algorithms for learning posteriors in deep and wide BNNs. On the other hand, by using normalizing flows to model weight priors they often can match target GP prior (see Sec. 4.2). Nevertheless, imposing different priors for each weight invalidates the theoretical assumptions regarding BNN convergence, rendering this approach merely a heuristic.
@@
Fitting only weight (and biases) priors is insufficient for matching BNN and GP priors. Hence, we propose learning parametric activations as well. Our proposed method transfers priors from GPs to BNNs by matching their function-space distributions using the closed-form 2-Wasserstein and by learning activations in addition to priors on weights. It achieves faithful function-space priors using just shallow BNNs. In contrast to Tran et al., (2022), it is backed by theoretical results ensuring that the learned BNNs asymptotically converge to GPs, and compared to Meronen et al., (2020), it can work with arbitrary kernels. Finally, it relies on simple weight priors, enabling efficient posterior inference.
@@
Within the proposed framework, we propose two novel extensions. First, by conditioning the BNN’s functional priors (weights and activation) on GP hyperparameters – implemented via hypernetworks – we facilitate efficient model selection, e.g., allow prior hyperparameters optimization. Second, building on the result of Meronen et al., (2021) that periodic activations induce stationarity, we propose trainable periodic activations and show how such activations can be learned in practice, ensuring global stationarity in BNNs.
@@
In the Appendix, the results presented in the paper are supplemented with a description of related work, a detailed overview of the experimental settings, and additional plots and numerical data.
Our code is made public at (github).
@@ 2 Preliminaries
Behavior akin to GPs has been identified in wide neural networks with a single hidden layer by Neal, (1996); Williams, (1996) and later generalized for deep architectures by Lee et al., (2017); Matthews et al., (2018) who considered stacking multiple such the wide layers.
@@ 3.1 Transfer of Priors from GP to BNN
In practice, however, we settle for approximate matching over finite index sets (Shi et al., , 2018; Sun et al., , 2019), where the densities over BNN and GP outputs should approximately align as $p_{nn}(f^{l}(X))\approx p_{gp}(f^{l}(X))$. $X\sim p_{X}$ ($X\in\mathcal{X}$) can be understood as sampling sets of inputs $\{x\}$ at which we observe the functions.
@@ 3.2 Differentiable Priors and Activations
The results presented in Sec. 2 hold for distributions with finite variances and light tails. We use the basic zero-centered factorized Gaussians with learned variances, e.g., $p(w|\sigma^{2}_{w})$ and $p(b|\sigma^{2}_{b})$, as prior distributions on model weights and biases. To enable gradient-based optimization, such as gradient propagation through weight and bias samples, we reparameterize the distributions using the reparameterization trick (Kingma and Welling, , 2014). Additionally, we propose to employ parametric and differentiable activations $\phi(\cdot|\eta)$ to enhance the network’s flexibility. This previously overlooked idea allows us to model complex functional priors within a single hidden layer of a BNN, even with simple prior distributions on weights and biases.
@@
We explore several models for $\phi$, including Rational (Pade) activations (Molina et al., , 2019), Piecewise Linear (PWL) activations [^1] https://pypi.org/project/torchpwl/, and activations implemented as a neural network with a single narrow (with only 5 neurons) hidden layer, using ReLU/SiLU own activations. These functions are computationally efficient and introduce desirable non-linear properties, striking a balance between efficiency and the capacity to model intricate patterns.
@@ 3.3 Periodic Activations For Stationary GPs
We address the local stationarity challenge by introducing trainable activations designed to exhibit periodic behavior. This approach is motivated by the result of Meronen et al., (2021), who showed that periodic activation functions induce stationarity in BNNs. Based on this theoretical result, we propose a practical solution. Our proposed activations
@@ 3.4 Conditional Priors and Activations
Researchers previously, when fitting BNN priors, were limited to upfront fixing hyperparameters (such as lengthscale) of target GPs or alternatively, they would employ workarounds like hierarchical kernels with hyperparameters sampled from hyperpriors (see, e.g., Tran et al., (2022)). Instead, we propose to condition the priors on weights, biases, and activations in BNNs on the hyperparameters of a GP kernel. By incorporating GP hyperparameters directly into BNNs, we bridge the equivalence gap between GPs and BNNs, finally enabling BNNs to fully replicate the behavior of GPs. In particular, integrating hyperparameters directly within BNNs allows for their optimization, for example, using marginal likelihood (evidence). This facilitates effective model selection, which is a significant novelty compared to previous approaches.
@@
The conditioning we implemented using hypernetworks (Ha et al., , 2017; Chauhan et al., , 2024), as $[\sigma,\eta]:=\text{hnet}(\gamma|\theta)$, where $\sigma$ and $\eta$ are the sets of parameters for the priors on weights and activations, respectively. Here, $\gamma$ represents the set of conditioning hyperparameters, which are transformed by the hypernetwork hnet. The hypernetwork has its own parameters $\theta$, which are now optimized in Eq. 2 as $\lambda=\{\theta\}$ instead of $\{\sigma,\eta\}$. For a fixed architecture of hnet, $\sigma$ and $\eta$ are now fully determined by $\gamma$ along with $\theta$.
@@ 3.5 Loss
However, unlike Tran et al., (2022), we propose to use the 2-Wasserstein metric, which for Multivariate Gaussians (applicable in the case of GPs and wide BNNs -- at least approximately) has a closed-form solution (Mallasto and Feragen, , 2017):
@@
where $\mu_{1/2}$, $\Sigma_{1/2}$ are respectively expectations and covariance matrices estimated for $p_{nn}$ and $p_{gp}$ from samples $\{f^{l}(X)\}$ (we used $512$ or $1024$ reparameterized samples) evaluated for inputs $X$. Not only we avoid the internal optimisation due to $\sup_{|\psi|}$, but additionally $D$ can be efficiently computed based on results by Buzuti and Thomaz, (2023). For experiments, we report numerical values of $D$ normalized by number of elements in $X$.
@@ 3.6 Output Structure
The optimization objective in Eq. 2 is defined over functions $f$. The functions are passed through likelihoods to form model outputs $y$. For regression tasks, a Gaussian likelihood is typically used; for binary classification, Bernoulli; and for multiclass classification, a Categorical likelihood with multivariate $f$ transformed via softmax. Note that the prior transfer is independent of a likelihood, meaning it does not rely on how the latent functions $f$ relate to the outputs $y$. Then, as explained in Sec. 3.5, the optimization can be performed efficiently between Gaussians. If these assumptions were not satisfied—e.g., if the desired priors are not expressible via a GP -- we can revert to the nested optimization for $|\Psi|_{L}$ as implied by Eq. 3. Overall, we explain our method in terms of GPs and for brevity, we denote the target functional priors by $p_{gp}$. However, it is important to note that the method is applicable to arbitrarily specified functional priors.
For completeness, we follow to briefly discuss also the transfer of priors from GPs to BNNs for multivariate/multi-class-output settings; however, such extensions fall outside of the scope of this paper. In particular, the multivariate models (Rasmussen and Williams, , 2005; Alvarez et al., , 2012) can be approached using various architectural strategies, depending on the nature of the dependencies among the output dimensions. One approach is to construct independent BNNs for each output dimension, stacking them side-by-side, where each BNN is pretrained separately with priors obtained from independent single-output GPs. This approach is appropriate when output dimensions are assumed to be independent, leading to an ensemble of independently trained BNNs. Alternatively, shared hidden layer BNN can be used, where a common hidden representation is shared across all output dimensions, thereby capturing potential dependencies between outputs. The shared hidden layer structure reflects the idea of a multi-output or multi-task GP, where dependencies among outputs can be modeled explicitly using coregionalization techniques (Goovaerts, , 1997; Bonilla et al., , 2008). The shared hidden layer BNN can thus be endowed with priors derived from these multi-output GPs, ensuring that both spatial and output correlations are incorporated into the BNN model.
@@ 4 Findings
Figure 1: Prior (a) and posterior (b) predictive distributions for a BNN with trained parameters priors and activations (ours; 4th column), and for Tran et al., (2022) approach with different prior realizations (Gaussian (3 hidden layers; 3rd column) and NF (2 hidden layers; 2nd). The first column illustrates the ground truth (GP). Numerical results complementing the figures we provide in the Appendix.
Table 1: Results for UCI regression task (Boston dataset) with prior transferred from a GP; comparison between the baseline (using a deep BNN; (Tran et al., , 2022)) and ours (single hidden layer BNN with varying widths). Periodic activation (Sec. 3.3) and an activation realized by a NN with SiLU own activation (default) were used.
@@
Metric $\downarrow$ - regularization + regularization - regularization + regularization - regularization + regularization ---
@@ 4.2 Do Expressive Function-space Priors Require Networks to be Deep?
Function-space priors require expressive models, which can be achieved either through a multi-layer neural network architecture or with a shallow BNN with more flexible (e.g., learnable) activations. (Tran et al., , 2022) explored the former, focusing on multi-layer networks. Instead of considering BNNs with wide layers -- where the theoretical results in Sec. 2 apply -- they postulated that a deep network can model a functional prior after tuning the network’s parameter priors to match a target functional prior.
@@
(Tran et al., , 2022) investigated several types of priors for BNN weights, including Gaussian, hierarchical, and Normalizing Flow (NF)-based priors. They argue that only the NF prior is flexible enough to capture complex distributions. While NFs (Rezende and Mohamed, , 2015) can model intricate priors across all weights, by assigning a distinct prior to each weight they violate the assumptions of the Central Limit Theorem (CLT), preventing the BNN from converging to a GP (see Sec. 2). Although (Tran et al., , 2022) achieved their best results using this heuristic NF prior, among the considered priors, the theoretical guarantees hold only for the Gaussian and hierarchical priors, and only in the context of wide BNNs.
@@
Beyond the lack of theoretical justification, using complex priors such as NFs or requiring deep networks introduces additional challenges in posterior inference. MCMC-based methods (Chen et al., , 2014; Del Moral et al., , 2006) become computationally expensive, while approximate inference techniques (Hoffman et al., , 2013; Ritter et al., , 2018) may lack expressiveness.
@@
In Fig. 2 we compare against our approach the behavior of their methods for varying numbers of layers and activations. Specifically, we evaluate posterior quality in a 1D regression task (Fig. 1), both on training data and in out-of-distribution (OOD) regions. The results show that multiple hidden layers are required for the Gaussian prior. In contrast, for the NF priors, increasing depth leads to worse performance in OOD regions -- illustrating an example of the heuristic’s failure. For both methods, we note high sensitivity to activation change. For example, NFs perform best for one or two layers, but only with RBF activation. For Swish or ReLU they fail to capture the data (high RMSE/NLL). For a more extensive analysis please see Sec. C in the Appendix.
Refer to captionRefer to captionRefer to captionRefer to caption
@@
Figure 2: Comparison of posterior predictive quality across the methods from (Tran et al., , 2022) for the problem from Fig. 1, considering different priors, activation functions, and numbers of hidden layers. We evaluate both out-of-distribution (in-between region) performance (left, Wasserstein divergence) and in-distribution performance (right, RMSE). For the Gaussian prior (top), the results align with the assumption by the authors -- adding more layers improves the GP approximation. Contrary, for the Normalizing Flow prior (bottom), the best performance is observed with one or two hidden layers, while deeper networks suffer from instability, highlighting failure of the heuristic. Our method (black line) consistently outperforms the baselines.
@@
For further evaluation, we compare our approach with Tran et al., (2022) on a range of regression tasks, including both synthetic and real-world datasets (e.g., the Boston dataset). The results are shown in Fig.1 and Tab.1, respectively. Additional discussion and results for several other regression tasks are provided in the Appendix. We closely follow the settings picked by Tran et al., (2022) and still observe that our method performs comparably or better than the baseline, while requiring only a single-hidden-layer BNN.

### v3

@@ 1 Introduction
Function-space priors for BNNs offer a better way of specifying beliefs on data modeled by the model. Unlike priors on model parameters (i.e. weights), which lack interpretability and are hard to specify, they ultimately lead to more intuitive and meaningful representation of prior knowledge [Sun et al., 2019]. Function-space priors can be conveniently specified in terms of GPs. Instead of specifying complex distributions over model parameters, GPs allow for defining priors over entire functions, through the choice of kernel capturing requirements such as smoothness or periodicity.
@@
Finding a GP that matches a wide BNN is straightforward. A classic result by Neal [1996], Williams [1996], later extended to deep neural networks by Lee et al. [2017], Matthews et al. [2018], shows that wide layers in neural networks behave a priori like GPs. This is achieved by identifying GP kernels matching the covariances of (B)NNs.
@@
In contrast, finding BNNs that exhibit the desired behavior of GPs is a notoriously difficult problem, with solutions or approximations existing for only a few GP kernels. For example, Meronen et al. [2020] recently found a solution for the popular Matérn kernel using the Wiener-Khinchin theorem. However, the proposed BNN activation assumes binary white noise priors on the model weights, which severely hinders posterior learning. Consequently, an approximate solution with Gaussian priors must be used, ultimately leading to suboptimal performance, as we demonstrate in Sec. 4.3.
@@
The challenge of imposing function-space a priori behavior on BNNs previously has been addressed using gradient-based optimization. In particular, Flam-Shepherd et al. [2017], Flam-Shepherd et al. [2018], and especially Tran et al. [2022] attempted to solve this by learning priors on parameters. However, to achieve sufficient fidelity with simple priors (Gaussian or hierarchical), their approach requires deep networks, which presents a challenge for posterior learning due to the lack of effective algorithms for learning posteriors in deep and wide BNNs. On the other hand, by using normalizing flows to model weight priors they often can match target GP prior (see Sec. 4.2). Nevertheless, imposing different priors for each weight invalidates the theoretical assumptions regarding BNN convergence, rendering this approach merely a heuristic.
@@
Fitting only weight (and biases) priors is insufficient for matching BNN and GP priors. Hence, we propose learning parametric activations as well. We draw inspiration from previous work demonstrating that function-space priors can be adjusted by altering activations [Neal, 1996] and also from research utilizing activations for improved uncertainty estimation [Morales-Alvarez et al., 2020, Postels et al., 2020]. In fact, the same GP-like behavior can often be realized by different combinations of parameter priors and activations, as seen in the BNN covariance formulation (Eq. 1) and learning both jointly offers a greater flexibility. In particular, our method (see Fig. 1) transfers priors from GPs to BNNs by matching their function-space distributions using the closed-form 2-Wasserstein loss and by learning activations in addition to priors on weights. It achieves faithful function-space priors using just shallow BNNs. In contrast to Tran et al. [2022], it is backed by theoretical results ensuring that the learned BNNs asymptotically converge to GPs, and compared to Meronen et al. [2020], it can work with arbitrary kernels. Finally, it relies on simple weight priors, enabling efficient posterior inference.
@@
Within the proposed framework, we propose two novel extensions. First, by conditioning the BNN’s functional priors (weights and activation) on GP hyperparameters – implemented via hypernetworks – we facilitate efficient model selection, e.g., allow prior hyperparameters optimization. Second, building on the result of Meronen et al. [2021] that periodic activations induce stationarity, we propose trainable periodic activations and show how such activations can be learned in practice, ensuring global stationarity in BNNs.
@@
In the Appendix, the results presented in the paper are supplemented with a description of related work, a discussion of computational advantages of our approach, a detailed overview of the experimental settings, and additional plots and numerical data.
Our code we made publicly available  [^1] https://github.com/gmum/bnn-functional-priors.
Figure 1: Schematic view of our approach.
@@ 2 Preliminaries
Behavior akin to GPs has been identified in wide neural networks with a single hidden layer by Neal [1996], Williams [1996] and later generalized for deep architectures by Lee et al. [2017], Matthews et al. [2018] who considered stacking multiple such the wide layers.
@@ 3.1 Transfer of Priors from GP to BNN
In practice, however, we settle for approximate matching over finite index sets [Shi et al., 2018, Sun et al., 2019], where the densities over BNN and GP outputs should approximately align as $p_{nn}(f^{l}(X))\approx p_{gp}(f^{l}(X))$. $X\sim p_{X}$ ($X\in\mathcal{X}$) can be understood as sampling sets of inputs $\{x\}$ at which we observe the functions.
@@ 3.2 Differentiable Priors and Activations
The results presented in Sec. 2 hold for distributions with finite variances and light tails. We use the basic zero-centered factorized Gaussians with learned variances, e.g., $p(w|\sigma^{2}_{w})$ and $p(b|\sigma^{2}_{b})$, as prior distributions on model weights and biases. To enable gradient-based optimization, such as gradient propagation through weight and bias samples, we reparameterize the distributions using the reparameterization trick [Kingma and Welling, 2014]. Additionally, we propose to employ parametric and differentiable activations $\phi(\cdot|\eta)$ to enhance the network’s flexibility. This previously overlooked idea allows us to model complex functional priors within a single hidden layer of a BNN, even with simple prior distributions on weights and biases.
@@
We explore several models for $\phi$, including Rational (Pade) activations [Molina et al., 2019], Piecewise Linear (PWL) activations [^2] https://pypi.org/project/torchpwl/, and activations implemented as a neural network with a single narrow (with only 5 neurons) hidden layer, using ReLU/SiLU own activations. These functions are computationally efficient and introduce desirable non-linear properties, striking a balance between efficiency and the capacity to model intricate patterns.
Our choice of a NN-based activation with a single hidden layer consisting of 5-10 neurons and using ReLU/SiLU activations was informed by empirical studies (see Fig. 2 and Tables 3 and 4 in Section B) comparing various learnable functions. These experiments, including ablations on the number of layers and width of the NN activations, show that this configuration offers a good balance of flexibility, fast convergence, and quality of prior fit. While deeper or wider NN activations can marginally improve fidelity in some cases, they also increase convergence times and can make optimization harder.
@@ 3.3 Periodic Activations For Stationary GPs
We address the local stationarity challenge by introducing trainable activations designed to exhibit periodic behavior. This approach is motivated by the result of Meronen et al. [2021], who showed that periodic activation functions induce stationarity in BNNs. Based on this theoretical result, we propose a practical solution. Our proposed activations
@@ 3.4 Conditional Priors and Activations
Researchers previously, when fitting BNN priors, were limited to upfront fixing hyperparameters (such as lengthscale) of target GPs or alternatively, they would employ workarounds like hierarchical kernels with hyperparameters sampled from hyperpriors (see, e.g., Tran et al. [2022]). Instead, we propose to condition the priors on weights, biases, and activations in BNNs on the hyperparameters of a GP kernel. By incorporating GP hyperparameters directly into BNNs, we bridge the equivalence gap between GPs and BNNs, finally enabling BNNs to fully replicate the behavior of GPs. In particular, integrating hyperparameters directly within BNNs allows for their optimization, for example, using marginal likelihood (evidence). This facilitates effective model selection, which is a significant novelty compared to previous approaches.
@@
The conditioning we implemented using hypernetworks [Ha et al., 2017, Chauhan et al., 2024], as $[\sigma,\eta]:=\text{hnet}(\gamma|\theta)$, where $\sigma$ and $\eta$ are the sets of parameters for the priors on weights and activations, respectively. Here, $\gamma$ represents the set of conditioning hyperparameters, which are transformed by the hypernetwork hnet. The hypernetwork has its own parameters $\theta$, which are now optimized in Eq. 2 as $\lambda=\{\theta\}$ instead of $\{\sigma,\eta\}$. For a fixed architecture of hnet, $\sigma$ and $\eta$ are now fully determined by $\gamma$ along with $\theta$.
@@ 3.5 Loss
However, unlike Tran et al. [2022], we propose to use the 2-Wasserstein metric, which for Multivariate Gaussians (applicable in the case of GPs and wide BNNs – at least approximately) has a closed-form solution [Mallasto and Feragen, 2017]:
@@
where $\mu_{1/2}$, $\Sigma_{1/2}$ are respectively expectations and covariance matrices estimated for $p_{nn}$ and $p_{gp}$ from samples $\{f^{l}(X)\}$ (we used $512$ or $1024$ reparameterized samples) evaluated for inputs $X$. Not only we avoid the internal optimisation due to $\sup_{|\psi|}$, but additionally $D$ can be efficiently computed based on results by Buzuti and Thomaz [2023]. For experiments, we report numerical values of $D$ normalized by number of elements in $X$.
@@ 3.6 Output Structure
The optimization objective in Eq. 2 is defined over functions $f$. The functions are passed through likelihoods to form model outputs $y$. For regression tasks, a Gaussian likelihood is typically used; for binary classification, Bernoulli; and for multiclass classification, a Categorical likelihood with multivariate $f$ transformed via softmax. Note that the prior transfer is independent of a likelihood, meaning it does not rely on how the latent functions $f$ relate to the outputs $y$. Then, as explained in Sec. 3.5, the optimization can be performed efficiently between Gaussians. If these assumptions were not satisfied—e.g., if the desired priors are not expressible via a GP – we can revert to the nested optimization for $|\Psi|_{L}$ as implied by Eq. 3. Overall, we explain our method in terms of GPs and for brevity, we denote the target functional priors by $p_{gp}$. However, it is important to note that the method is applicable to arbitrarily specified functional priors.
For completeness, we follow to briefly discuss also the transfer of priors from GPs to BNNs for multivariate/multi-class-output settings; however, such extensions fall outside of the scope of this paper. In particular, the multivariate models [Rasmussen and Williams, 2005, Alvarez et al., 2012] can be approached using various architectural strategies, depending on the nature of the dependencies among the output dimensions. One approach is to construct independent BNNs for each output dimension, stacking them side-by-side, where each BNN is pretrained separately with priors obtained from independent single-output GPs. This approach is appropriate when output dimensions are assumed to be independent, leading to an ensemble of independently trained BNNs. Alternatively, shared hidden layer BNN can be used, where a common hidden representation is shared across all output dimensions, thereby capturing potential dependencies between outputs. The shared hidden layer structure reflects the idea of a multi-output or multi-task GP, where dependencies among outputs can be modeled explicitly using coregionalization techniques [Goovaerts, 1997, Bonilla et al., 2008]. The shared hidden layer BNN can thus be endowed with priors derived from these multi-output GPs, ensuring that both spatial and output correlations are incorporated into the BNN model.
@@ 4 Findings
Table 1: Results for UCI regression task (Boston dataset) with prior transferred from a GP; comparison between the baseline (using a deep BNN; [Tran et al., 2022]) and ours (single hidden layer BNN with varying widths). Periodic activation (Sec. 3.3) and an activation realized by a NN with SiLU own activation (default) were used.
@@
Metric $\downarrow$ - regularization + regularization - regularization + regularization - regularization + regularization
@@
input space dimensionality = 1
input space dimensionality = 16
Figure 2: Quality of matching BNNs to the prior of a GP with a Matérn kernel ($\nu=5/2$, $\ell=1$) for 1D (top) and 16D inputs (bottom). We evaluate models with trained parameter priors (denoted by w), activations (denoted by a), and both (denoted by a+w). Each label specifies whether a fixed activation (e.g., ReLU) or a specific activation model (e.g., Rational) was employed. Gaussian parameter priors were used by default. If not trained, we set variances to $1$, and for the hidden layer, we normalized the variance by its width. The label w: Matérn refers to a BNN with the closed-form (fixed) activation as derived by Meronen et al. [2020].
@@ 4.1 Does Learning Activations Improve Learning of Function-space Priors?
Figure 3: Prior (a) and posterior (b) predictive distributions for a BNN with trained parameters priors and activations (ours; 4th column), and for Tran et al. [2022] approach with different prior realizations (Gaussian (3 hidden layers; 3rd column) and NF (2 hidden layers; 2nd). The first column illustrates the ground truth (GP). Numerical results complementing the figures we provide in the Appendix.
@@ 4.2 Do Expressive Function-space Priors Require Networks to be Deep?
Function-space priors require expressive models, which can be achieved either through a multi-layer neural network architecture or with a shallow BNN with more flexible (e.g., learnable) activations. [Tran et al., 2022] explored the former, focusing on multi-layer networks. Instead of considering BNNs with wide layers – where the theoretical results in Sec. 2 apply – they postulated that a deep network can model a functional prior after tuning the network’s parameter priors to match a target functional prior.
@@
[Tran et al., 2022] investigated several types of priors for BNN weights, including Gaussian, hierarchical, and Normalizing Flow (NF)-based priors. They argue that only the NF prior is flexible enough to capture complex distributions. While NFs [Rezende and Mohamed, 2015] can model intricate priors across all weights, by assigning a distinct prior to each weight they violate the assumptions of the Central Limit Theorem (CLT), preventing the BNN from converging to a GP (see Sec. 2). Although [Tran et al., 2022] achieved their best results using this heuristic NF prior, among the considered priors, the theoretical guarantees hold only for the Gaussian and hierarchical priors, and only in the context of wide BNNs.
@@
Beyond the lack of theoretical justification, using complex priors such as NFs or requiring deep networks introduces additional challenges in posterior inference. MCMC-based methods [Chen et al., 2014, Del Moral et al., 2006] become computationally expensive, while approximate inference techniques [Hoffman et al., 2013, Ritter et al., 2018] may lack expressiveness.
@@
In Fig. 4 we compare our approach against the behavior of their methods for varying numbers of layers and activations. Specifically, we evaluate posterior in a 1D regression task (Fig. 3), both on training data and in out-of-distribution (OOD) regions. The results show that multiple hidden layers are required for the Gaussian prior. In contrast, for the NF priors, increasing depth leads to worse performance in OOD regions – illustrating an example of the heuristic’s failure. For both methods, we note high sensitivity to activation change. NFs perform best for one or two layers, but only with RBF activation. For Swish/ReLU they fail to capture the data (high RMSE/NLL). Please see also Sec. E.
@@
Figure 4: Comparison of posterior predictive quality across the methods from [Tran et al., 2022] for the problem from Fig. 3, considering different priors, activation functions, and numbers of hidden layers. We evaluate both out-of-distribution (in-between region) performance (left, Wasserstein divergence) and in-distribution performance (right, RMSE). For the Gaussian prior (top), the results align with the assumption by the authors – adding more layers improves the GP approximation. Contrary, for the Normalizing Flow prior (bottom), the best performance is observed with one or two hidden layers, while deeper networks suffer from instability, highlighting failure of the heuristic. Our method (black line) consistently outperforms the baselines.
@@
For further evaluation, we compare our approach with Tran et al. [2022] on a range of regression tasks, including both synthetic and real-world datasets (e.g., the Boston dataset). The results are shown in Fig.3 and Tab.1, respectively. Additional discussion and results for several other regression tasks are provided in the Appendix. We closely follow the settings picked by Tran et al. [2022] and still observe that our method performs comparably or better than the baseline, while requiring only a single-hidden-layer BNN.